Find the volume of the solid generated by revolving the region bounded by the curves about the y axis y=x^3 , x=0 , y=1


1
Expert's answer
2021-12-24T08:52:13-0500
"y=x^3=>x=\sqrt[3]{y}"

The solid lies between "y =0" and "y= 1," its volume is

"V=\displaystyle\int_{0}^1\pi(\sqrt[3]{y})^2dy=\pi[\dfrac{3}{5}y^{5/3}]\begin{matrix}
 1 \\
 0
\end{matrix}=0.6\pi({units}^3)"